OpenStudy (anonymous):

for every 3 laps that Jackie swims, Ronnie swims 4 laps. During practice, they swam a combined total of 63 laps. how many laps did Ronnie swim during practice?

OpenStudy (anonymous):

36 laps..!

OpenStudy (agent0smith):

Call the number of laps Jackie swims J and Ronnie R. They swam a total of 63 laps: J + R = 63 "for every 3 laps that Jackie swims, Ronnie swims 4 laps" This one is a bit tougher to explain, but notice he swims 4/3 times as many laps as her \[\large \frac{ 4}{3 }J = R\]

OpenStudy (agent0smith):

\[\large J + R = 63\] \[\large \frac{ 4}{3 }J = R\] You can solve this using substitution.

OpenStudy (anonymous):

let jackie swim 3x laps..and ronnie swin 4x laps .. 3x+4x=63 x=9 so ronnie swam 4x that is 4 multiplied by 9 as 36 !